# Is it possible to describe an electron by the evolution of a single point?

### v2

@@ Abstract
The answer to the title-question is affirmative. The analysis of the geometry of continuous and differentiable curves in three-dimensional Euclidean space suggests that the point represents the location of the center of charge of the electron, satisfies a system of ordinary differential equations of fourth order, and moves at the speed of light. The center of mass of the electron is a different point and will be determined by the evolution of the center of charge. It is the relative motion of the center of charge around the center of mass that gives rise to the spin and magnetic properties. The invariance of the mass and the absolute value of the spin for the center of mass observer imply that in the interaction of the electron with an external electromagnetic field the particle has to radiate.
@@ 1 Introduction
Let us consider the possibility of describing a classical spinning electron by the evolution of a single point $\bi r$} in three-dimensional Euclidean space. From the physical point of view it represents a mechanical system of three degrees of freedom. If we achieve this, it means that all observables related to the structure of the electron, like energy, linear momentum, spin, magnetic moment, electric dipole moment, etc., can be expressed in terms of the point $\bi r$} and its subsequent time derivatives.
@@ 1.1 Geometry
If we take in (2) the next order derivative $d^{4}{\bi r}/ds^{4}$, taking into account (4) and (3) it can be expressed as a linear combination of the three unit vectors where the coefficients are functions of the four derivatives. These three unit vectors are expressed in (1) in terms of the derivatives of the point $\bi r$} up to the third order, and therefore we obtain a relationship between the derivatives of $\bi r$} up to order four. This means that in the Frenet-Serret formalism the most general differential equation satisfied by a continuous and differentiable geometric curve in three dimensional space is a system of ordinary differential equations of fourth order where the independent variable is the arc length $s$.
@@ 1.2 Physics
Let us go one step forward. Instead of doing Geometry let us make Physics by introducing clocks. Let us consider that we describe the electron by the time evolution in three-dimensional space of a point $\bi r$}. In the above geometrical analysis it will mean that there exists a continuous and differentiable function $s(t)$ between the arc length $s$ and the elapsed time $t$, that will represent from now on the parameter of the curve ${\bi r}(t)\equiv{\bi r}(s(t))$. Let us consider that an inertial observer describes the free motion of this electron. Because in general, all derivatives are different from zero, the motion of the point is accelerated and therefore this point ${\bi r}(t)$ cannot represent the location of the center of mass (CM) of the electron. It can represent perhaps the location of another property of the electron.
If we are able to obtain a Lagrangian description of this system of only three degrees of freedom, the free Lagrangian $L_{0}$ has to depend up to the second time derivative of the point $\bi r$}, to obtain from the Euler-Lagrange equations a system of differential equations of fourth order. If this electron interacts there will be an interaction Lagrangian $L_{I}$, function of the interacting parameter, the charge, the point $\bi r$} and some of its time derivatives and of external fields defined at point $\bi r$}. This is telling us that the point $\bi r$} represents the point where the external fields are evaluated. The external force is evaluated at this point, i.e., it represents the center of charge (CC) of the electron. This point has to be a different point than the possible CM, because in the free motion the CM must describe a straight line with no acceleration.
@@
Let us consider that in an inertial reference frame the motion of point $\bi r$}, although accelerated, represents a free motion. This means that for this inertial observer the displacement of the point at any time $t$, $ds(t)$, must be independent of the time. Otherwise if at two different times $t_{1}$ and $t_{2}$, $ds(t_{1})\neq ds(t_{2})$ this means that something different is happening at time $t_{2}$ than at time $t_{1}$, that produces a different displacement, contradictory with the idea that the motion is free. For this inertial observer the free motion of the center of charge of the electron is moving at a velocity of constant absolute value. If this description has to be independent of the inertial observer we arrive to the conclusion that we can never find an inertial observer at rest with respect to this point.
@@
The general motion of the point $\bi r$} has curvature and torsion. If it is moving at a velocity of constant absolute value for some inertial observer, the velocity is no longer of constant absolute value for another inertial observer moving at a constant velocity $\bi v$} with respect to the first one. If we add a constant vector $\bi v$} to a changing velocity vector does not produce a vector of a constant absolute value. There is nevertheless a unique solution if the velocities do not add according to the nonrelativistic linear addition. The velocity of the CC of an electron described by a single point must necessarily be an unreachable velocity of constant absolute value for all inertial observers. The velocity of the point $\bi r$} must transform according to the relativistic addition formula and it must be the speed of light. This is consistent with Dirac’s analysis of the electron such that the point $\bi r$} of Dirac’s spinor $\psi(t,{\bi r})$, where the external fields are evaluated, is in fact moving at the speed of light for all inertial observers. Dirac’s spinor represents the probability amplitude of the electron around the CC.
@@ 1.3 Invariance arguments
Let us assume that an inertial observer describes the evolution of the electron by the trajectory of a point ${\bi r}(t)$ in three-dimensional space. If we make an arbitrary Poincaré transformation to any other inertial observer what we get is the family of all trajectories ${\bi r}(t,g)$ of the point for all inertial observers, where $t$ is the time coordinate in each frame. This family is parameterized by the 10 parameters that characterize the arbitrary Poincaré transformation $g$. All these trajectories satisfy a Poincaré invariant differential equation obtained from the above family ${\bi r}(t,g)$ and its subsequent time derivatives, by eliminating the 10 parameters in terms of the different derivatives. To eliminate 10 parameters we need to reach up to the fourth-order derivative to have sufficient equations, and therefore we find again that the trajectory of the point for any inertial observer satisfies a system of Poincaré invariant ordinary fourth order differential equations. This dynamical equation is Poincaré invariant by construction because is independent of all group parameters.
The point is not the center of mass, and in the free motion, if it represents the center of charge it is accelerated but cannot radiate because energy is conserved. A particle with CC and CM as different points can never radiate if the CM is not accelerated. A free electron never radiates. The Lorentz-Dirac theory of radiation considers that both points are the same, but this is an unphysical constraint. The center of mass represents the location of the center of inertia of the particle while the CC represents the location of the center of interaction. Inertia and interaction are different physical properties of matter. Theoretical physics is oblidged to analyze physical systems where these two points could be different.
@@ 2 General description
The previous introduction is contained basically in the Preamble of the Lecture Notes [1]. However, from a historical point of view, we arrived at this description of the electron by a different path. In 1985 [2] we analyzed the Lagrangian description of a mechanical system such that the boundary variables manifold that define the initial and final states of its variational evolution were considered to be a homogeneous space of the Galilei group. With this statement many models of non-relativistic spinning particles were described. The relativistic version of the above models were obtained in 1989 [3] by assuming that the boundary variables manifold of the variational formalism were homogeneous spaces of the Poincaré group. In that article we also assumed that the velocity of the point $\bi r$} could be the speed of light, as suggested by Dirac’s analysis. Different models of relativistic spinning particles were found and also a spinning model of a classical photon described as a point and a comoving cartesian frame that rotates along the direction of the velocity. The spin of the photon has the same direction than the angular velocity, but it is an invariant mechanical property. The angular velocity and thus, the frequency of the photon, transform according to the Doppler effect transformation while the absolute value of the spin remains invariant.
@@
The idea underlying this definition of a classical elementary particle is based on Wigner’s definition of an elementary particle in quantum formalism. According to Wigner the Hilbert space of the states of an elementary particle is a representation space of a unitary irreducible representation of the Poincaré group [4]. Irreducible means that the complete Hilbert space can be generated from any state. By selecting any arbitrary state, we can obtain all Poincaré transformations of this arbitrary state. Now let us produce all finite linear combinations of these new states and finally the closure of these states by including all limits of Cauchy sequences of this kind of vectors. We have generated the whole Hilbert space. This is the meaning that the representation is irreducible. This is a group theoretical definition of elementary particle.
@@
The quantum mechanical theory of elementary particles is also based on three fundamental principles. In the classical description, the Lagrangian formalism determines the path of the system between the initial point $x_{1}$ to the final point $x_{2}$. Because we are not going to measure the path followed by the system we assume Feynman’s hypothesis that all paths between these two points are allowed with the same probability. We have to substitute the Variational Principle by Feynman’s Quantization Principle such that the Hilbert space that describes the quantum mechanical states is such that any wave function is a complex, squared integrable function defined on the boundary variables manifold of the corresponding classical Lagrangian description [5]-[6].
@@
The formalism can be extended to any kinematical group, Galilei, Poincaré, Weyl, Conformal or De Sitter groups, or any group we consider that describes the space-time symmetries of our mechanical system. If the rotation group is a subgroup of this general kinematical group we always obtain that the elementary particles described in this formalism have spin. If for instance, we select as the boundary variables manifold the space-time, i.e. the states are defined by the variables time $t$ and the position of a point $\bi r$}, $x\equiv(t,{\bi r})$, this manifold is in fact a homogeneous space of the Galilei or Poincaré group, and therefore it can be used to describe an elementary particle: The spinless point particle. The angular momentum of this particle with respect to the point $\bi r$} is zero. But in Nature there are no spinless elementary particles. We need to include some more variables that transform under rotations, but we are restricted to a set, larger than space-time, that belongs to a homogeneous space of the kinematical group. In that case, the angular momentum of the particle with respect to the point $\bi r$} is different from zero and the particle has spin.
@@
In the quantum version, Dirac’s spinor is a squared integrable complex function of the variables $\psi(t,{\bi r},{\bi u},\boldsymbol{\alpha})$. The constraint $u=c$ implies that it is a function of the non-compact variables $(t,{\bi r})$ and of the following five compact variables: Two are the zenithal angle $\beta$ and azimuthal angle $\lambda$} of the velocity $\bi u$} of constant value $u=c$ and the other three angles that describe the orientation of the attached local cartesian frame. If we represent this local frame as a rotation matrix it is characterized by the rotated angle $\alpha$ and the zenithal angle $\theta$ and azimuthal angle $\phi$} of the unit vector along of the rotation axis. The spin operators are linear differential operators with respect to these 5 compact variables and commute with the Hamiltonian $H=i\hbar\partial/\partial t$ and the linear momentum operator ${\bi p}=-i\hbar\nabla$.
@@ 2.1 The Classical Dirac Particle
In addition to the previous references, we have published an article [9] where we analyze in detail what we call the classical Dirac particle, i.e., the relativistic model obtained by the general formalism that satisfies Dirac’s equation when quantized. All observables can be expressed in terms of the point $\bi r$} and its subsequent time derivatives. In particular, the center of mass (CM) $\bi q$} is defined with simultaneity with the CC, at every inertial frame, and it moves with a velocity ${\bi v}=d{\bi q}/dt$, such that $v<c$. The CC velocity ${\bi u}=d{\bi r}/dt$ always satisfies $u=c$. In terms of the CM velocity it is found that the energy and linear momentum can be written as
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. In the derivation of these dynamical equations we have made no use of a possible radiation reaction force. Equation (7) is just $d{\bi p}/dt={\bi F}$, after taking the time derivative of ${\bi p}=\gamma(v)m{\bi v}$, and leaving on the left hand side the CM acceleration $d{\bi v}/dt$. The equation (8) is equation (6) when writing on the left hand side the CC acceleration $\bi a$}.
@@
Now the electron has two characteristic points $\bi q$} and $\bi r$} and therefore two different angular momenta can be defined with respect to both points. By the Noether analysis it is found that the angular momentum with respect to the CC at time $t$ takes the form:
@@
The angular momentum with respect to the CM is defined as
@@
These observables represent the spin of the electron $\bi S$} and ${\bi S}_{CM}$ with respect to the CC and CM, respectively at the same time $t$. In the center of mass frame (${\bi q}={\bi v}=0$) and both spins take the same value. They satisfy two different dynamical equations. In the free case $d{\bi S}/dt={\bi p}\times{\bi u}$ and $d{\bi S}_{CM}/dt=0$. The CM spin is conserved but the CC spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It corresponds to the classical spin observable equivalent to the quantum mechanical Dirac’s spin. In the center of mass frame and for a free electron, the spin is constant and takes the value
@@
In the center of mass frame the separation $R_{0}=|{\bi r}-{\bi q}|$ is constant and when quantizing the model takes the value $R_{0}=\hbar/2mc$, and the internal motion has an angular velocity $\omega_{0}=2mc^{2}/\hbar$, twice the frequency postulated by De Broglie. The value of the spin is $S(0)=\hbar/2$.
@@
In natural units, the angular velocity is related to the Darboux vector $\boldsymbol{\omega}=\tau{\bi t}+\kappa{\bi b}$, so that the torsion is the component $\omega_{u}$ along the CC velocity and the curvature $\omega_{p}$ is along the binormal. The two components of the angular velocity of the comoving rotating frame correspond to the two intrinsic geometric properties of the trajectory of the CC.

### v3

@@ Abstract
The answer to the title-question is affirmative. The analysis of the geometry of continuous and differentiable curves in three-dimensional Euclidean space suggests that the point represents the location of the center of charge of the electron, satisfies a system of ordinary differential equations of fourth order, and moves at the speed of light. The center of mass of the electron is a different point and will be determined by the evolution of the center of charge. It is the relative motion of the center of charge around the center of mass that gives rise to the spin and magnetic properties. The invariance of the mass and the absolute value of the spin for the center of mass observer imply that in the interaction of the electron with an external electromagnetic field the particle has to radiate. The analysis of a Poincaré invariant interaction of two electrons implies that the only relevant parameter that characterizes the interaction in a natural system of units, is the fine structure constant. The fundamentals of General Relativity have to be revisited.
@@ 1 Introduction
Let us consider the possibility of describing a classical spinning electron by the evolution of a single point $\bi r$} in three-dimensional Euclidean space. From the physical point of view, it represents a mechanical system of three degrees of freedom. If we achieve this, it means that all observables related to the structure of the electron, like energy, linear momentum, spin, magnetic moment, electric dipole moment, etc., can be expressed in terms of the point $\bi r$} and its subsequent time derivatives.
@@ 1.1 Geometry
If we take in (2) the next order derivative $d^{4}{\bi r}/ds^{4}$, taking into account (4) and (3) it can be expressed as a linear combination of the three-unit vectors where the coefficients are functions of the four derivatives. These three-unit vectors are expressed in (1) in terms of the derivatives of the point $\bi r$} up to the third order, and therefore we obtain a relationship between the derivatives of $\bi r$} up to order four. This means that in the Frenet-Serret formalism the most general differential equation satisfied by a continuous and differentiable geometric curve in three-dimensional space is a system of ordinary differential equations of fourth order where the independent variable is the arc length $s$.
@@ 1.2 Physics
Let us go one step forward. Instead of doing Geometry let us make Physics by introducing clocks. Let us consider that we describe the electron by the time evolution in three-dimensional space of a point $\bi r$}. In the above geometrical analysis, it will mean that there exists a continuous and differentiable function $s(t)$ between the arc length $s$ and the elapsed time $t$, that will represent from now on the parameter of the curve ${\bi r}(t)\equiv{\bi r}(s(t))$. Let us consider that an inertial observer describes the free motion of this electron in terms of the time coordinate in its frame. Because in general, all derivatives are different from zero, the motion of the point is accelerated and therefore this point ${\bi r}(t)$ cannot represent the location of the center of mass (CM) of the electron. It can represent perhaps the location of another property of the electron.
If we are able to obtain a Lagrangian description of this system of only three degrees of freedom, the free Lagrangian $L_{0}$ has to depend up to the second time derivative of the point $\bi r$}, to obtain from the Euler-Lagrange equations a system of differential equations of fourth order. If this electron interacts there will be an interaction Lagrangian $L_{I}$, function of the interacting parameter, the charge, the point $\bi r$} and some of its time derivatives and of external fields defined at point $\bi r$}. This is telling us that the point $\bi r$} represents the point where the external fields are evaluated. The external force is evaluated at this point, i.e., $\bi r$} represents the center of charge (CC) of the electron. This point has to be a different point than the possible CM, because in the free motion the CM must describe a straight line with no acceleration.
@@
Let us consider that in an inertial reference frame the motion of point $\bi r$}, although accelerated, represents a free motion. This means that for this inertial observer the displacement of the point at any time $t$, $ds(t)$, must be independent of the time. Otherwise, if at two different times $t_{1}$ and $t_{2}$, $ds(t_{1})\neq ds(t_{2})$ this means that something different is happening at time $t_{2}$ than at time $t_{1}$, that produces a different displacement, contradictory with the idea that the motion is free. For this inertial observer the free motion of the center of charge of the electron is moving at a velocity of constant absolute value. If this description has to be independent of the inertial observer, we arrive to the conclusion that we can never find an inertial observer at rest with respect to this point.
@@
The general motion of the point $\bi r$} has curvature and torsion. If it is moving at a velocity of constant absolute value for some inertial observer, the velocity is no longer of constant absolute value for another inertial observer moving at a constant velocity $\bi v$} with respect to the first one, if we assume that velocities are added as in the non-relativistic formalism. If we add a constant vector $\bi v$} to a changing velocity vector does not produce a vector of a constant absolute value. There is nevertheless a unique solution if the velocities do not add according to the nonrelativistic linear addition. The velocity of the CC of an electron described by a single point must necessarily be an unreachable velocity of constant absolute value for all inertial observers. The velocity of the point $\bi r$} must transform according to the relativistic addition formula and it must be the speed of light. This is consistent with Dirac’s analysis of the electron such that the point $\bi r$} of Dirac’s spinor $\psi(t,{\bi r})$, where the external fields are evaluated, is in fact moving at the speed of light for all inertial observers. Dirac’s spinor represents the probability amplitude of the electron around the CC.
@@ 1.3 Invariance arguments
From another point of view, let us assume that an inertial observer describes the evolution of the electron by the trajectory of a point ${\bi r}(t)$ in three-dimensional space. If we make any arbitrary Poincaré transformation $g$ to any other inertial observer, what we get is the family of all trajectories ${\bi r}(t,g)$ of the point for all inertial observers, where $t$ is the time coordinate in each frame. This family is parameterized by the 10 parameters that characterize the arbitrary Poincaré transformation $g$. All these trajectories satisfy a Poincaré invariant differential equation obtained from the above family ${\bi r}(t,g)$ and its subsequent time derivatives, by eliminating the 10 parameters in terms of the different derivatives. To eliminate 10 parameters, we need to reach up to the fourth-order derivative to have sufficient equations, and therefore we find again that the trajectory of the point for any inertial observer satisfies a system of Poincaré invariant ordinary fourth order differential equations. This dynamical equation is Poincaré invariant by construction because is independent of all group parameters.
The point is not the center of mass, and in the free motion, if it represents the center of charge, it is accelerated but cannot radiate because energy is conserved. A particle with CC and CM as different points can never radiate if the CM is not accelerated. A free electron never radiates. The Lorentz-Dirac theory of radiation considers that both points are the same, but this is an unphysical constraint. The center of mass represents the location of the center of inertia of the particle while the CC represents the location of the center of interaction. Inertia and interaction are different physical properties of matter. Theoretical physics is obliged to analyze physical systems in which these two points could be different.
@@ 2 General description
The previous introduction is contained basically in the Preamble of the Lecture Notes [1]. However, from a historical point of view, we arrived at this description of the electron by a different path. In 1985 [2] we analyzed the Lagrangian description of a mechanical system such that the boundary variables manifold that define the initial and final states of its variational evolution was considered to be a homogeneous space of the Galilei group. With this statement many models of non-relativistic spinning particles were described. The relativistic version of the above models was obtained in 1989 [3] by assuming that the boundary variables manifold of the variational formalism were homogeneous spaces of the Poincaré group. In that article we also assumed that the velocity of the point $\bi r$} could be the speed of light, as suggested by Dirac’s analysis. Different models of relativistic spinning particles were found and also a spinning model of a classical photon described as a point and a comoving cartesian frame that rotates along the direction of the velocity. The spin of the photon has the same direction than the angular velocity, but it is an invariant mechanical property. The angular velocity and thus, the frequency of the photon, transform according to the Doppler effect transformation while the absolute value of the spin remains invariant.
@@
The idea underlying this definition of a classical elementary particle is based on Wigner’s definition of an elementary particle in quantum formalism. According to Wigner the Hilbert space of the states of an elementary particle is a representation space of a unitary irreducible representation of the Poincaré group [4]. Irreducible means that the complete Hilbert space can be generated from any state. By selecting any arbitrary state, we can obtain all Poincaré transformations of this arbitrary state. Now let us produce all finite linear combinations of these new states and finally the closure of these states by including all limits of Cauchy sequences of this kind of vectors. We have generated the whole Hilbert space. This is the meaning that the representation is irreducible. This is a group theoretical definition of elementary particle in the quantum formalism.
@@
The quantum mechanical theory of elementary particles is also based on three fundamental principles. In the classical description, the Lagrangian formalism determines the path of the system between the initial point $x_{1}$ to the final point $x_{2}$. Because we are not going to measure the path followed by the system, we assume Feynman’s hypothesis that all paths between these two points are allowed with the same probability. We have to substitute the Variational Principle by Feynman’s Quantization Principle such that the Hilbert space that describes the quantum mechanical states is such that any wave function is a complex, squared integrable function defined on the boundary variables manifold of the corresponding classical Lagrangian description [5]-[6].
@@
The formalism can be extended to any kinematical group, Galilei, Poincaré, Weyl, Conformal or De Sitter groups, or any group we consider that describes the space-time symmetries of our mechanical system. If the rotation group is a subgroup of this general kinematical group, we always obtain that the elementary particles described in this formalism have spin. If for instance, we select as the boundary variables manifold the space-time, i.e. the states are defined by the variables time $t$ and the position of a point $\bi r$}, $x\equiv(t,{\bi r})$, this manifold is in fact a homogeneous space of the Galilei or Poincaré group, and therefore it can be used to describe an elementary particle: The spinless point particle. The angular momentum of this particle with respect to the point $\bi r$} is zero. But in Nature there are no spinless elementary particles. We need to include some more variables that transform under rotations, but we are restricted to a set, larger than space-time, that belongs to a homogeneous space of the kinematical group. In that case, the angular momentum of the particle with respect to the point $\bi r$} is different from zero and the particle has spin.
@@
In the quantum version, Dirac’s spinor is a squared integrable complex function of the variables $\psi(t,{\bi r},{\bi u},\boldsymbol{\alpha})$. The constraint $u=c$ implies that $\psi$ is a function of the non-compact variables $(t,{\bi r})$ and of the following five compact variables: Two are the zenithal angle $\beta$ and azimuthal angle $\lambda$} of the velocity $\bi u$} of constant value $u=c$ and the other three, the angles that describe the orientation of the attached local cartesian frame. If we represent this local frame as a rotation matrix it is characterized by the rotated angle $\alpha$ and the zenithal angle $\theta$ and azimuthal angle $\phi$} of the unit vector along of the rotation axis. The spin operators are linear differential operators with respect to these 5 compact variables and commute with the Hamiltonian $H=i\hbar\partial/\partial t$ and the linear momentum operator ${\bi p}=-i\hbar\nabla$. By appropiately selecting the complete commuting set of operators, Dirac’s spinor can be written is separate variables as
$\psi(t,{\bi r},{\bi u},\boldsymbol{\alpha})=\sum_{i=1}^{i=4}\Psi_{i}(t,{\bi r})\chi_{i}(\beta,\lambda;\alpha,\theta,\phi),$
where the $\Psi_{i}(t,{\bi r})$ describe the space-time part and the four components $\chi_{i}(\beta,\lambda;\alpha,\theta,\phi)$ the internal or spin part associated to the spin operators. The different representations of Dirac’s equation, either Pauli-Dirac, Weyl or Majorana, etc., depend on how we describe the internal orientation operators [5].
@@ 2.1 The Classical Dirac Particle
In addition to the previous references, we have published an article [9] where we analyze in detail what we call the classical Dirac particle, i.e., the relativistic model obtained by the general formalism that satisfies Dirac’s equation when quantized. All observables can be expressed in terms of the point $\bi r$} and its subsequent time derivatives. In particular, the center of mass (CM) $\bi q$} is defined with simultaneity with the CC, at every inertial frame, and it moves with a velocity ${\bi v}=d{\bi q}/dt$, such that $v<c$. The CC velocity ${\bi u}=d{\bi r}/dt$ always satisfies $u=c$. In terms of the CM velocity, it is found that the energy and linear momentum can be written as
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force defined at the CC, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. In the derivation of these dynamical equations, we have made no use of a possible radiation reaction force. Equation (7) is just $d{\bi p}/dt={\bi F}$, after taking the time derivative of ${\bi p}=\gamma(v)m{\bi v}$, and leaving on the left-hand side the CM acceleration $d{\bi v}/dt$. The equation (8) is equation (6) when writing on the left-hand side the CC acceleration $\bi a$}.
@@
Now the electron has two characteristic points $\bi q$} and $\bi r$} and therefore two different angular momenta can be defined with respect to both points. By the Noether analysis it is found that the angular momentum with respect to the CC $\bi r$}, at time $t$ takes the form:
@@
The angular momentum at time $t$, with respect to the CM $\bi q$}, is defined as
@@
These observables represent the spin of the electron $\bi S$} and ${\bi S}_{CM}$ with respect to the CC and CM, respectively at the same time $t$. In the center of mass frame (${\bi q}={\bi v}=0$) and both spins take the same value. They satisfy two different dynamical equations. In the free case $d{\bi S}/dt={\bi p}\times{\bi u}$ and $d{\bi S}_{CM}/dt=0$. The CM spin is conserved but the CC spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It corresponds to the classical spin observable equivalent to the quantum mechanical Dirac’s spin operator. In the center of mass frame and for a free electron, the spin is constant and takes the value
@@
In the center of mass frame, the separation $R_{0}=|{\bi r}-{\bi q}|$ is constant and when quantizing the model takes the value $R_{0}=\hbar/2mc$, and the internal motion has an angular velocity $\omega_{0}=2mc^{2}/\hbar$, twice the frequency postulated by De Broglie. The value of the spin is $S(0)=\hbar/2$.
@@
In natural units, the angular velocity is related to the Darboux vector $\boldsymbol{\omega}=\tau{\bi t}+\kappa{\bi b}$, so that the torsion is the component $\omega_{u}$ along the CC velocity and the curvature $\omega_{p}$ is along the binormal. The two components of the angular velocity of the comoving rotating frame correspond to the two intrinsic geometric properties of the trajectory of the CC in terms of time derivatives instead of derivatives with respect to the arc length $s$.